# Reconstructing Hand-Held Objects in 3D from Images and Videos

### v3

@@ 7.5 Ablation Studies
MCC Baseline. Because our model is built on top of MCC [89], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction. For MCC, we use the released model that was trained on the CO3Dv2 dataset [67] and apply an existing monocular depth estimator [63] to obtain depth maps (as MCC takes an RGB-D image as input). Table 3 compares 3D metrics between MCC and MCC-HO computed on the MOW test set. We find that using MCC-HO, which was trained on the specific task of hand-held object reconstruction, leads to performance improvements on test images.
@@
MCC [89] 0.10 0.21 31.6
@@
RAR Success Rate. In order to quantify the success rate of RAR (e.g., whether a generated 3D object matches the input image), we use GPT-4(V) and Genie to automatically generate objects for all 512 images in the MOW dataset. Genie is able to generate an object 82% of the time, and otherwise returned an error. The results are included in the supplementary materials.
Video Alignment Components. We aim to understand the contributions of each cost term in our Viterbi algorithm, e.g., MCC-HO and DINOv2. See Table 4. In order to best assess the quality of rigid alignment, the ICP postprocess described in Section 7.2 is not applied. The results demonstrate that both terms contribute to the efficacy of our Viterbi algorithm; in practice, we observe that using both 2D and 3D cues is particularly important for objects that lie outside the training distribution of MCC-HO (DexYCB consists of commonplace YCB objects).
Viterbi Cost Terms F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC-HO 0.26 0.52 4.18
DINOv2 0.23 0.45 6.38
MCC-HO + DINOv2 0.27 0.52 4.10
Table 4: Analysis of how each of our Viterbi cost terms contributes to the final temporal alignment on the DexYCB test dataset.
@@ 8 Conclusion
In this paper, we present a novel paradigm for reconstructing hand-held objects that combines the respective benefits of model-free and model-based prediction. Our transformer-based model, MCC-HO, is trained to predict hand-object geometry given a single RGB image and estimated 3D hand. Experiments conducted using the DexYCB, MOW, and HOI4D datasets all demonstrate that MCC-HO achieves state-of-the-art performance in hand-held object reconstruction. We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets. These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH videos [71].

### v4

@@ 7.5 Ablation Studies
MCC Baseline. Because our model is built on top of MCC [wu2023multiview], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction. For MCC, we use the released model that was trained on the CO3Dv2 dataset [reizenstein21co3d] and apply an existing monocular depth estimator [Ranftl2021] to obtain depth maps (as MCC takes an RGB-D image as input). Table 3 compares 3D metrics between MCC and MCC-HO computed on the MOW test set. We find that using MCC-HO, which was trained on the specific task of hand-held object reconstruction, leads to performance improvements on test images.
@@
MCC [wu2023multiview] 0.10 0.21 31.6
MCC-HO 0.15 0.31 15.2
Table 3: Analysis of MCC-HO performance gains over our object reconstruction baseline, MCC, using the MOW test dataset.
Effectiveness of Input Hand. Table 4 analyzes the effectiveness of hand pose prediction input by evaluating MCC-HO with and without the input hand mesh. As we expect, the model performs significantly worse without the input hand geometry. The experimental results further corroborate the hypothesis that grounding the 3D object prediction with respect to the hand coordinate system significantly constrains the reconstruction problem.
Model Parameters F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC [wu2023multiview] 0.10 0.21 31.6
MCC-HO (w/o hand) 0.03 0.12 71.6
@@
RAR Reconstruction Improvements. RAR leverages text-to-3D generative models to synthesize realistic geometry, and we quantitatively evaluate whether using RAR in conjunction with MCC-HO leads to improved results. Table 5 demonstrates that RAR (with Genie objects) can indeed help improve the object reconstruction quality of monocular reconstruction methods such as MCC-HO.
Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC-HO 0.15 0.31 15.2
MCC-HO + RAR 0.16 0.44 9.41
Table 5: Analysis of how RAR improves object reconstruction. Median Chamfer Distance (cm2) and F-score (5mm, 10mm) metrics are computed on the MOW test dataset.
@@ 8 Conclusion
In this paper, we present a novel paradigm for reconstructing hand-held objects that combines the respective benefits of model-free and model-based prediction. Our transformer-based model, MCC-HO, is trained to predict hand-object geometry given a single RGB image and estimated 3D hand. Experiments conducted using the DexYCB, MOW, and HOI4D datasets all demonstrate that MCC-HO achieves state-of-the-art performance in hand-held object reconstruction. We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets. These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH videos [shan2020understanding].